# Fractal transference principles for subsets of $\mathbb{N}^{d}$ of positive density

### v1

@@ Abstract
We establish a multidimensional fractal transference principle for digit-restricted sets associated with subsets of $\mathbb{N}^{d}$, extending the one-dimensional framework of Nakajima–Takahasi [24]. We develop general Hausdorff-dimension tools via the singular value potential $\phi^{s}(\mathbf{a})$ and the multivariate Dirichlet series $\zeta_{S}(\bm{\sigma})=\sum_{\mathbf{a}\in S}\prod_{j=1}^{d}a_{j}^{-\sigma_{j}}$. Let $s_{\ast}:=\inf\{s>0:\sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a})<\infty\}$ and $\Lambda_{S}:=\inf\{\sigma_{1}+\cdots+\sigma_{d}:\zeta_{S}(\bm{\sigma})<\infty\}$. We obtain $\dim_{H}(\mathcal{E}_{S})\leq s_{\ast}$, where $\mathcal{E}_{S}\subset(0,1)^{d}$ denotes the set of points whose continued-fraction digit vectors lie in $S$ and whose coordinates escape (i.e. $a_{n}(x_{j})\to\infty$ for each $j$), and $s_{\ast}=\tfrac{1}{2}\Lambda_{S}$ for uniformly KK–balanced $S$. In particular, if $S\subset\mathbb{N}^{d}$ has positive upper (or upper Banach) density then $\dim_{H}(\mathcal{E}_{S})=d/2$. On the combinatorial side, the transference principle ensures that translation-invariant configurations forced at positive density, including multidimensional Szemerédi patterns, persist inside the induced fractal digit sets.
@@ 1 Introduction
These examples emphasize that digit restrictions encode arithmetic information in a geometric way. In this paper, however, we work with the coordinatewise continued fraction expansion and impose digit constraints through subsets of $\mathbb{N}^{d}$.
@@ 1.1 Coordinatewise continued fractions and digit sets
The study of $E_{1}$ was suggested by P. Erdős; see Ramharter [25, p. 11]. Ramharter [25] proved that $\dim_{H}(E_{1})=\tfrac{1}{2}$; see also Jordan–Rams [16]. For comparison, Good’s landmark theorem [10, Theorem 1] asserts that the escaping set
@@
Similar product phenomena occur in other symbolic codings. For example, Cheng–Zhang [5, Theorem 1.4] study exact-approximation sets arising from $\beta$-expansions and obtain a sharp formula for the Hausdorff dimension of their Cartesian products when $d\geq 2$. At the level of critical Hausdorff measure, even very classical self-similar products can be delicate: Guo–Jones [11] provide improved upper and lower bounds for the critical Hausdorff measure of Cartesian products of the middle-third Cantor set, extending what was previously known in dimension two.
@@ 1.2 Fractal transference principles on $\mathbb{N}^{d}$
Our first main theorem is an $\mathbb{N}^{d}$-analogue of the one-dimensional transference principle of Nakajima–Takahasi [24, Theorem 1.1]. Informally, it asserts that one can choose Hausdorff-typical points in the escaping, vector-injective regime $\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ so that their digit sets reflect the density of a prescribed subset $S\subset\mathbb{N}^{d}$.
Before stating our results, we recall the one-dimensional theorem of Nakajima–Takahasi [24, Theorem 1.1]. In dimension $d=1$, Nakajima–Takahasi [24, Theorem 1.1] proved that if $S\subset\mathbb{N}$ has $\overline{d}(S)>0$ (resp. $d^{\ast}(S)>0$), then there exist sets $E_{S}$ (resp. $F_{S}$) with
@@
When $d=1$, (1)–(4) recover the fractal transference principle of Nakajima–Takahasi [24, Theorem 1.1]. Our proof follows the same surgery–and–elimination paradigm, but several points are genuinely high–dimensional: one must enforce coordinatewise divergence ($\min_{1\leq i\leq d}a_{n}(x_{i})\to\infty$) and vector–injectivity ($\mathbf{a}_{n}(\mathbf{x})\neq\mathbf{a}_{m}(\mathbf{x})$ for $n\neq m$), and the dimension analysis of the model families is anisotropic, governed by dd–dimensional cylinder geometry rather than one–dimensional lengths.
@@ 1.3 Relative fractal transference on $\mathbb{N}^{d}$
We also develop a relative version of Nakajima–Takahasi [24, Theorem 1.4], where a sparse ambient set $S\subset\mathbb{N}^{d}$ plays the role of a background set. For sharp upper bounds it is convenient to control anisotropy: for $K\geq 1$ let
@@
Before stating our higher-dimensional result, we recall the one-dimensional theorem of Nakajima–Takahasi [24, Theorem 1.4]. In dimension $d=1$, they they showed that if $S\subset\mathbb{N}$ is infinite with polynomial density of exponent $\alpha\geq 1$ and $A\subset S$ satisfies $\overline{d}(A\mid S)>0$, then there exists a set $E_{S,A}\subset\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}}$ with
@@
Let $S\subset\mathcal{C}_{K}\subset\mathbb{N}^{d}$ be infinite with polynomial density of exponent $\alpha\geq 1$, and let $A\subset S$ satisfy $\overline{d}(A\mid S)>0$. Then
$\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}})=\frac{d}{2\alpha},$
@@
$\dim_{H}\Bigl\{\mathbf{x}\in\mathcal{E}_{S}^{\mathrm{vec}}:\ \overline{d}\bigl(D(\mathbf{x})\cap A\mid S\bigr)=\overline{d}(A\mid S)\Bigr\}=\dim_{H}(\mathcal{E}_{S}^{\mathrm{vec}}).$
@@
The uniform balance hypothesis excludes extremely anisotropic digit vectors and places the polynomial-growth assumption in a regime where our covering argument applies cleanly and yields sharp upper bounds; without it, polynomial density alone need not be sufficient for our method.
  1. (i)
    When $d=1$ this hypothesis is automatic, so Theorem 1.4 recovers the one-dimensional result of Nakajima–Takahasi [24, Theorem 1.4] without any additional geometric restriction.
  2. (ii)
    In higher dimensions it is natural to ask how far the balance condition can be relaxed. For instance, in strongly anisotropic situations where admissible digits grow at very different rates across coordinates, it is unclear how the distribution of $S$ among directions should affect the dimension of $\mathcal{E}_{S}^{\mathrm{vec}}$, and whether one should expect product-type behaviour reflecting only partial independence between directions.
@@ 2.1 Some facts on continued fractions
    For any two irrational numbers $x_{1},x_{2}\in(0,1)$ with the same first $n$ partial quotients $a_{1},\dots,a_{n}$ but $a_{n+1}(x_{1})\neq a_{n+1}(x_{2})$, there exists an absolute constant $c>0$ such that
@@
Both statements are standard consequences of the elementary geometry of continued fraction cylinders and can be verified by a direct computation. See, for instance, [10, 13, 24]. ∎
@@
$|I(a_{1},\dots,a_{n})|=\frac{1}{q_{n}(q_{n}+q_{n-1})}\leq\frac{1}{q_{n}^{2}}.$
In particular, $|I(a_{1},\dots,a_{n})|\asymp q_{n}^{-2}$ with absolute constants.
@@
Let $\mathbf{x}_{1},\mathbf{x}_{2}\in(0,1)^{d}$ such that $\mathbf{a}_{i}(\mathbf{x}_{1})=\mathbf{a}_{i}(\mathbf{x}_{2})$ for $1\leq i\leq n$, but $\mathbf{a}_{n+1}(\mathbf{x}_{1})\neq\mathbf{a}_{n+1}(\mathbf{x}_{2})$. Then there exists a constant $C>0$ depending only on $d$ such that
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C\cdot\max_{j\in\mathcal{D}}\left(|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{|a_{n+1}^{(j)}(\mathbf{x}_{1})-a_{n+1}^{(j)}(\mathbf{x}_{2})|}{a_{n+1}^{(j)}(\mathbf{x}_{1})a_{n+1}^{(j)}(\mathbf{x}_{2})}\right),$
where $\mathcal{D}=\{1\leq j\leq d:a_{n+1}^{(j)}(\mathbf{x}_{1})\neq a_{n+1}^{(j)}(\mathbf{x}_{2})\}$.
@@
By the product structure from Lemma 2.4, the cylinder set $I(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})$ decomposes as a Cartesian product of one-dimensional cylinder sets:
$I(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})=I^{(1)}_{\omega}\times I^{(2)}_{\omega}\times\cdots\times I^{(d)}_{\omega}$,
where each $I^{(j)}_{\omega}$ is the one-dimensional cylinder determined by $(a_{1}^{(j)},\ldots,a_{n}^{(j)})$.
For each coordinate $j\in\mathcal{D}$, the one-dimensional continued fractions $x_{1,j}$ and $x_{2,j}$ have the same first $n$ partial quotients $a_{1}^{(j)},\ldots,a_{n}^{(j)}$ but differ at the $(n+1)$-th partial quotient. By Lemma 2.1(ii), there exists an absolute constant $c_{0}>0$ such that for each $j\in\mathcal{D}$,
$|x_{1,j}-x_{2,j}|\;\geq\;\frac{c_{0}}{q_{n,j}^{2}}\,\frac{\bigl|a_{n+1}(x_{1,j})-a_{n+1}(x_{2,j})\bigr|}{a_{n+1}(x_{1,j})\,a_{n+1}(x_{2,j})},$
where $q_{n,j}$ is the denominator of the $n$-th convergent associated with $(a_{1}^{(j)},\ldots,a_{n}^{(j)})$. Since $|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\leq q_{n,j}^{-2}$ by Lemma 2.2, we may rewrite the lower bound as
$|x_{1,j}-x_{2,j}|\;\geq\;c\,|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{\bigl|a_{n+1}(x_{1,j})-a_{n+1}(x_{2,j})\bigr|}{a_{n+1}(x_{1,j})\,a_{n+1}(x_{2,j})},$
for some absolute constant $c>0$ (e.g. $c=c_{0}$).
Since $\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}=\max_{1\leq j\leq d}|x_{1,j}-x_{2,j}|$, we have:
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq\max_{j\in\mathcal{D}}|x_{1,j}-x_{2,j}|\geq c\cdot\max_{j\in\mathcal{D}}\left(|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{|a_{n+1}^{(j)}(\mathbf{x}_{1})-a_{n+1}^{(j)}(\mathbf{x}_{2})|}{a_{n+1}^{(j)}(\mathbf{x}_{1})a_{n+1}^{(j)}(\mathbf{x}_{2})}\right).$
Taking $C=c$ completes the proof. Note that the constant $c$ here is the same absolute constant from the one-dimensional case and does not depend on $d$. ∎
The separation properties of these cylinder sets are inherited from the corresponding one-dimensional continued fraction geometry.
@@
In the later construction of the seed (root) set, the $n$-th level pieces are chosen inside dd–dimensional continued–fraction cylinders. To verify the vanishing mesh condition in Definition 4.1, it is convenient to have a uniform estimate showing that one–dimensional cylinders shrink exponentially in the word length, with constants independent of the digits. A bound of the form $|I(a_{1},\dots,a_{n})|\ll\prod_{k=1}^{n}a_{k}^{-2}$ is not uniform when some digits are small, and therefore it is not suitable for deducing $\max_{Q\in\mathcal{Q}_{n}}\operatorname{diam}(Q)\to 0$ without additional growth assumptions. The following lemma is a basic property of regular continued fractions; we include a complete proof for convenience.
@@
Let $p_{n}/q_{n}$ be the $n$-th convergent associated with $(a_{1},\dots,a_{n})$, with $q_{-1}=0$, $q_{0}=1$, and
$q_{n}=a_{n}q_{n-1}+q_{n-2}\qquad(n\geq 1).$
It is standard that the cylinder $I(a_{1},\dots,a_{n})$ is the open interval whose endpoints are the two rationals
$\frac{p_{n}}{q_{n}}\quad\text{and}\quad\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}$
(up to order, depending on the parity of $n$). Using the determinant identity $p_{n}q_{n-1}-p_{n-1}q_{n}=(-1)^{n-1}$, we obtain the exact length formula
$|I(a_{1},\dots,a_{n})|=\left|\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}-\frac{p_{n}}{q_{n}}\right|=\frac{1}{q_{n}(q_{n}+q_{n-1})}\leq\frac{1}{q_{n}^{2}}.$
Since $a_{n}\geq 1$ for all $n$, the recursion yields $q_{n}\geq q_{n-1}+q_{n-2}$, and hence $q_{n}\geq F_{n+1}$ for all $n\geq 0$, where $(F_{m})$ is the Fibonacci sequence. By Binet’s formula,
$F_{m}=\frac{\varphi^{m}-\psi^{m}}{\sqrt{5}},\qquad\psi=\frac{1-\sqrt{5}}{2},$
so $F_{m}\geq c\,\varphi^{m}$ for all $m\geq 1$ with some absolute constant $c>0$. Combining these estimates gives
$|I(a_{1},\dots,a_{n})|\leq\frac{1}{F_{n+1}^{2}}\leq C\,\varphi^{-2n},$
which proves the claim (taking $\vartheta=\varphi^{-2}$ and adjusting $C$ if needed). ∎

### v2

@@ Abstract
We establish a multidimensional fractal transference principle for digit-restricted sets associated with subsets of $\mathbb{N}^{d}$, extending the one-dimensional framework of Nakajima–Takahasi (Adv. Math., 2025). We develop general Hausdorff-dimension tools via the singular value potential $\phi^{s}(\mathbf{a})$ and the multivariate Dirichlet series $\zeta_{S}(\bm{\sigma})=\sum_{\mathbf{a}\in S}\prod_{j=1}^{d}a_{j}^{-\sigma_{j}}$. Let $s_{\ast}:=\inf\{s>0:\sum_{\mathbf{a}\in S}\phi^{s}(\mathbf{a})<\infty\}$ and $\Lambda_{S}:=\inf\{\sigma_{1}+\cdots+\sigma_{d}:\zeta_{S}(\bm{\sigma})<\infty\}$. We obtain $\dim_{H}(\mathcal{E}_{S})\leq s_{\ast}$, where $\mathcal{E}_{S}\subset(0,1)^{d}$ denotes the set of points whose continued-fraction digit vectors lie in $S$ and whose coordinates escape (i.e. $a_{n}(x_{j})\to\infty$ for each $j$), and $s_{\ast}=\tfrac{1}{2}\Lambda_{S}$ for uniformly KK–balanced $S$. In particular, if $S\subset\mathbb{N}^{d}$ has positive upper (or upper Banach) density then $\dim_{H}(\mathcal{E}_{S})=d/2$. On the combinatorial side, the transference principle ensures that translation-invariant configurations forced at positive density, including multidimensional Szemerédi patterns, persist inside the induced fractal digit sets.
@@ 1 Introduction
These examples emphasize that digit restrictions encode arithmetic information in a geometric way. In this paper, however, we work with the coordinatewise continued fraction expansion and impose digit constraints through subsets of $\mathbb{N}^{d}$. More generally, once different directions contract at different rates, sharp Hausdorff-dimension bounds typically require multi-scale coverings and can depend on how the allowed pieces are distributed among directions, even in diagonal linear models; see McMullen [33] and the survey of Fraser [11], and also Barański [1]. Our goal is to combine this geometric viewpoint with density input from additive combinatorics, via a higher-dimensional fractal transference principle.
@@ 1.1 Coordinatewise continued fractions and digit sets
The study of $E_{1}$ was suggested by P. Erdős; see Ramharter [36, p. 11]. Ramharter [36] proved that $\dim_{H}(E_{1})=\tfrac{1}{2}$; see also Jordan–Rams [24]. For comparison, Good’s landmark theorem [14, Theorem 1] asserts that the escaping set
@@
Similar product phenomena occur in other symbolic codings. For example, Cheng–Zhang [6, Theorem 1.4] study exact-approximation sets arising from $\beta$-expansions and obtain a sharp formula for the Hausdorff dimension of their Cartesian products when $d\geq 2$. At the level of critical Hausdorff measure, even very classical self-similar products can be delicate: Guo–Jones [15] provide improved upper and lower bounds for the critical Hausdorff measure of Cartesian products of the middle-third Cantor set, extending what was previously known in dimension two.
@@ 1.2 Fractal transference principles on $\mathbb{N}^{d}$
Our first main theorem is an $\mathbb{N}^{d}$-analogue of the one-dimensional transference principle of Nakajima–Takahasi [34, Theorem 1.1]. Informally, it asserts that one can choose Hausdorff-typical points in the escaping, vector-injective regime $\mathcal{E}_{\mathbb{N}^{d}}^{\mathrm{vec}}$ so that their digit sets reflect the density of a prescribed subset $S\subset\mathbb{N}^{d}$.
Before stating our results, we recall the one-dimensional theorem of Nakajima–Takahasi [34, Theorem 1.1]. In dimension $d=1$, Nakajima–Takahasi [34, Theorem 1.1] proved that if $S\subset\mathbb{N}$ has $\overline{d}(S)>0$ (resp. $d^{\ast}(S)>0$), then there exist sets $E_{S}$ (resp. $F_{S}$) with
@@
When $d=1$, (1)–(4) recover the fractal transference principle of Nakajima–Takahasi [34, Theorem 1.1]. Our proof follows the same surgery–and–elimination paradigm, but several points are genuinely high–dimensional: one must enforce coordinatewise divergence ($\min_{1\leq i\leq d}a_{n}(x_{i})\to\infty$) and vector–injectivity ($\mathbf{a}_{n}(\mathbf{x})\neq\mathbf{a}_{m}(\mathbf{x})$ for $n\neq m$), and the dimension analysis of the model families is anisotropic, governed by dd–dimensional cylinder geometry rather than one–dimensional lengths.
@@ 1.3 Relative fractal transference on $\mathbb{N}^{d}$
We also develop a relative version of Nakajima–Takahasi [34, Theorem 1.4], where a sparse ambient set $S\subset\mathbb{N}^{d}$ plays the role of a background set. For sharp upper bounds it is convenient to control anisotropy: for $K\geq 1$ let
@@
Before stating our higher-dimensional result, we recall the one-dimensional theorem of Nakajima–Takahasi [34, Theorem 1.4]. In dimension $d=1$, they they showed that if $S\subset\mathbb{N}$ is infinite with polynomial density of exponent $\alpha\geq 1$ and $A\subset S$ satisfies $\overline{d}(A\mid S)>0$, then there exists a set $E_{S,A}\subset\mathcal{E}_{\mathbb{N}}^{\mathrm{vec}}$ with
@@
Let $S\subset\mathcal{C}_{K}\subset\mathbb{N}^{d}$ be infinite with polynomial density of exponent $\alpha\geq 1$. For any $A\subset S$ with $\overline{d}(A\mid S)>0$, there exists a set $E_{S,A}\subset\mathcal{E}_{S}^{\mathrm{vec}}$ such that
$\dim_{H}E_{S,A}=\dim_{H}\bigl(\mathcal{E}_{S}^{\mathrm{vec}}\bigr)=\frac{d}{2\alpha},$
and
$\overline{d}\Bigl(\bigcup_{n\geq 1}\ \bigcap_{\mathbf{x}\in E_{S,A}}\{\mathbf{a}_{n}(\mathbf{x})\}\cap A\ \Bigm|\ S\Bigr)=\overline{d}(A\mid S),$
@@
$\dim_{H}\Bigl\{\mathbf{x}\in\mathcal{E}_{S}^{\mathrm{vec}}:\ \overline{d}\bigl(D(\mathbf{x})\cap A\mid S\bigr)=\overline{d}(A\mid S)\Bigr\}=\dim_{H}\bigl(\mathcal{E}_{S}^{\mathrm{vec}}\bigr).$
@@
When $d=1$ the balance condition is automatic, so Theorem 1.4 recovers Nakajima–Takahasi [34, Theorem 1.4].
The following examples are included only to indicate why we impose the balance condition: without $S\subset\mathcal{C}_{K}$, different coordinates may contract at different rates, and dimension can become sensitive to directional distribution rather than a single counting exponent.
Remark 1.6.
This type of obstruction already appears in linear diagonal self-affine models; see McMullen [33], where the Hausdorff dimension is computed for planar self-affine carpets (often called Bedford–McMullen carpets). Fix integers $n>m\geq 2$ and a digit set $D\subset\{0,\ldots,n-1\}\times\{0,\ldots,m-1\}$. For $(i,j)\in D$ define $f_{i,j}(x,y)=((x+i)/n,(y+j)/m)$ and let $K_{D}\subset[0,1]^{2}$ be the attractor $K_{D}=\bigcup_{(i,j)\in D}f_{i,j}(K_{D})$. When $n\neq m$ the contraction rates differ, and $\dim_{H}K_{D}$ is not determined by #D\#D alone: it depends on the row counts $t_{j}=\#\{\,i:(i,j)\in D\,\}$; see [33, §1, Theorem]. In particular, two carpets with the same #D\#D can have different Hausdorff dimensions because their digits are distributed differently.
Although our sets come from continued fractions (Gauss map) rather than an affine IFS, the same mechanism can occur: without $S\subset\mathcal{C}_{K}$, different coordinates may contract at incompatible rates at the same depth, so a single growth exponent such as $|S\cap Q_{N}|$ may not capture the cylinder geometry sharply enough for our one-scale covering and seed bounds to match.
See also Lalley–Gatzouras [26] and the affinity-dimension framework of Falconer [9] and Hueter–Lalley [21] for related non-conformal phenomena. For a survey perspective, see Fraser [11], where it is stressed that even for these diagonal linear models the Hausdorff-dimension upper bound typically uses delicate multi-scale coverings and depends on how the allowed rectangles are distributed.
Remark 1.7.
A further comparison is Barański [1], who studies planar rectangle-like constructions extending the carpet setting. One has nested compact sets $\Delta_{\varepsilon_{1},\ldots,\varepsilon_{n}}\subset\mathbb{R}^{2}$ ($\varepsilon_{i}\in\{1,\ldots,M\}$) with $\Delta_{\varepsilon_{1},\ldots,\varepsilon_{n},\varepsilon_{n+1}}\subset\Delta_{\varepsilon_{1},\ldots,\varepsilon_{n}}$ and $\operatorname{diam}(\Delta_{\varepsilon_{1},\ldots,\varepsilon_{n}})\to 0$, and the limit set
$\Lambda=\bigcap_{n\geq 1}\ \bigcup_{\varepsilon_{1},\ldots,\varepsilon_{n}}\Delta_{\varepsilon_{1},\ldots,\varepsilon_{n}}.$
In the anisotropic case the dimension formulas already require optimisation: $\dim_{H}\Lambda=\max_{q\in S}g(q)$ [1, Theorem A], where $g$ is an explicit continuous entropy/Lyapunov-type function depending on the directional distribution of symbols, and $\dim_{B}\Lambda=\max(D_{A},D_{B})$ [1, Theorem B], where $D_{A},D_{B}$ are the two box-dimension candidates given by pressure-type equations for the horizontal/vertical scales (hence depending on $(a_{i}),(b_{j})$ and the digit pattern). In particular, $\dim_{H}\Lambda$ and $\dim_{B}\Lambda$ may differ.
Although our cylinders come from the non-linear Gauss map rather than an affine IFS, the same issue is relevant: without balance, different coordinates can produce different effective scales at the same depth, so a single counting exponent (e.g. for $|S\cap Q_{N}|$) may be insufficient for sharp one-scale covering bounds.
See also Peres–Shmerkin [35] and Hochman–Shmerkin [20] for resonance/overlap effects in product-type settings, and Jenkinson–Pollicott [23] and Hensley [17] for transfer-operator approaches in continued-fraction dimension problems.
@@ 2.1 Some facts on continued fractions
    For any two irrational numbers $x_{1},x_{2}\in(0,1)$ with the same first $n$ partial quotients $a_{1},\dots,a_{n}$ but
    $|a_{n+1}(x_{1})-a_{n+1}(x_{2})|\geq 2,$
    there exists an absolute constant $c>0$ such that
@@
Both statements are standard consequences of the elementary geometry of continued fraction cylinders and can be verified by a direct computation. See, for instance, [8, 14, 19, 34]. ∎
@@
$\frac{1}{2q_{n}^{2}}\leq|I(a_{1},\dots,a_{n})|=\frac{1}{q_{n}(q_{n}+q_{n-1})}\leq\frac{1}{q_{n}^{2}}.$
@@
Let $\mathbf{x}_{1},\mathbf{x}_{2}\in(0,1)^{d}$ satisfy $\mathbf{a}_{i}(\mathbf{x}_{1})=\mathbf{a}_{i}(\mathbf{x}_{2})$ for $1\leq i\leq n$ and
$\bigl\|\mathbf{a}_{n+1}(\mathbf{x}_{1})-\mathbf{a}_{n+1}(\mathbf{x}_{2})\bigr\|_{\infty}\geq 2.$
Let $\mathcal{D}:=\{1\leq j\leq d:\ |a_{n+1}^{(j)}(\mathbf{x}_{1})-a_{n+1}^{(j)}(\mathbf{x}_{2})|\geq 2\}$. Then there exists a constant $C>0$ depending only on $d$ such that
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C\max_{j\in\mathcal{D}}\left(|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{|a_{n+1}^{(j)}(\mathbf{x}_{1})-a_{n+1}^{(j)}(\mathbf{x}_{2})|}{a_{n+1}^{(j)}(\mathbf{x}_{1})\,a_{n+1}^{(j)}(\mathbf{x}_{2})}\right).$
@@
By Lemma 2.4,
$I(\mathbf{a}_{1},\ldots,\mathbf{a}_{n})=\prod_{j=1}^{d}I(a_{1}^{(j)},\ldots,a_{n}^{(j)}).$
Fix $j\in\mathcal{D}$. Then $x_{1,j}$ and $x_{2,j}$ have the same first $n$ partial quotients $(a_{1}^{(j)},\ldots,a_{n}^{(j)})$ and $|a_{n+1}(x_{1,j})-a_{n+1}(x_{2,j})|\geq 2$. By Lemma 2.1(ii),
$|x_{1,j}-x_{2,j}|\geq\frac{c_{0}}{q_{n,j}^{2}}\,\frac{|a_{n+1}(x_{1,j})-a_{n+1}(x_{2,j})|}{a_{n+1}(x_{1,j})\,a_{n+1}(x_{2,j})},$
where $q_{n,j}$ is the denominator of the $n$-th convergent associated with $(a_{1}^{(j)},\ldots,a_{n}^{(j)})$. Since $|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\leq q_{n,j}^{-2}$ by Lemma 2.2, we obtain
$|x_{1,j}-x_{2,j}|\geq c_{0}\,|I(a_{1}^{(j)},\ldots,a_{n}^{(j)})|\cdot\frac{|a_{n+1}(x_{1,j})-a_{n+1}(x_{2,j})|}{a_{n+1}(x_{1,j})\,a_{n+1}(x_{2,j})}.$
Taking the maximum over $j\in\mathcal{D}$ and using $\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}=\max_{1\leq j\leq d}|x_{1,j}-x_{2,j}|$ gives the claim (with $C=c_{0}$). ∎
The next lemma gives a uniform quantitative separation estimate between sibling cylinders, and will be used to verify the separation condition in Definition 4.1.
@@
To verify the vanishing mesh condition in Definition 4.1, we record the following classical lemma.
@@
Let $p_{n}/q_{n}$ be the $n$-th convergent of $(a_{1},\dots,a_{n})$. The endpoint formula gives
$|I(a_{1},\dots,a_{n})|=\frac{1}{q_{n}(q_{n}+q_{n-1})}\leq\frac{1}{q_{n}^{2}}.$
Since $a_{k}\geq 1$, the recursion $q_{n}=a_{n}q_{n-1}+q_{n-2}$ implies $q_{n}\geq q_{n-1}+q_{n-2}$, hence $q_{n}\geq F_{n+1}$ for all $n$, where $(F_{n})$ is the Fibonacci sequence. Moreover, $F_{n+1}\geq\varphi^{\,n-1}$ for $n\geq 1$ (by induction using $\varphi^{2}=\varphi+1$). Therefore
$|I(a_{1},\dots,a_{n})|\leq F_{n+1}^{-2}\leq\varphi^{2}\,\varphi^{-2n}.$
Thus one may take $\vartheta=\varphi^{-2}$ and $C=\varphi^{2}$. ∎